f(x)=3x^-1 for x=2/3
Answer by clarkp(14)

f(x)=3x^-1 for x=2/3
f(2/3)= 3(2/3)^-1
=6/3^-1
=2^-1
=therefore the answer is 1/2.

clark,, cheerr!!!